OpenStudy (anonymous):

find the indefinite integral: ∫(x)/(1+e^(-x^(2)))dx I know this question can be solved by "integration by table", but ...can we solve it with General power rule? thanks

OpenStudy (anonymous):

\[\int\limits \frac{ x }{ 1+e ^{-x ^{2}} }\]

OpenStudy (shubhamsrg):

simplify the expression first . (remove negative exponent)

sam (.sam.):

u sub? try u=x^2 -du/2=xdx \[-\frac{1}{2} \int\limits \frac{du}{1+e^u} \]

OpenStudy (shubhamsrg):

well, if you simply first, then you'll find a better term to substitute for, try it.

OpenStudy (shubhamsrg):

simplify* first

OpenStudy (experimentx):

\[ \frac{1}{1+e^u} = 1 - \frac{e^u}{1+e^u}\] Also assuming u>0 \[ \frac{1}{1+e^u} = e^{-u}\sum_{n=0}^\infty (-1)^{n}e^{-nu}\] if you integrate it .. you should get log series

OpenStudy (anonymous):

my method is ... first \[(\frac{ x }{ 1+e ^{-x ^{2}} })(\frac{ e ^{x ^{2}} }{ e ^{x ^{2}} })\] and then become... \[\int\frac{ xe ^{x ^{2}} }{ e ^{x ^{2}}+1 }dx\] \[{ u=e ^{x ^{2}}+1 }\] ... but the ans looks not correct... did I do something wrong?

OpenStudy (anonymous):

tabluar integration x+1|e^(x^2+2x) 1 |e^(x^2+2x)/(2x+2) 0 |e^(x^2+2x)/(2x+2)^2 y=(x+1)(e^(x^2+2x)/(2x+2))-(e^(x^2+2x)…

OpenStudy (anonymous):

Let I =\[\int\limits \frac{ x }{ 1+e ^{-x ^{2}} }dx\] \[substitute x ^{2}=t,diff. 2xdx=dt,xdx=\frac{ dt }{2 }\] \[I=\frac{ 1 }{ 2 }\int\limits \frac{ dt }{1+e ^{-t} }\] multiply numerator & denominator by e^t \[I=\frac{ 1 }{ 2 }\int\limits \frac{ e ^{t} }{e ^{t}+1 }dt ,I=\frac{ 1 }{ 2 } \log \left( e ^{t}+1 \right)+c\] substitute the value of t & get the answer.